You buy 50 feet of fence material, you plan to plant a garden within the fence,
and you are going to get some fertilizer. One bag of fertilizer covers 500 square feet, you decide that you want to make your garden so that one bag of fertilizer is just enough.
what length do you make the first side of your fence?

Answer by richwmiller(17219)

circular garden
2* pi* r=50
pi*r^2=500
not possible
rectangular garden
2L+2W=50
L*W=500
not possible
must be some other shape